Determine coefficients a and b such that p(x)=x2+ax+b satisfies p(1)=−6 and p′(1)=3.

a=


b=
1
Expert's answer
2020-11-10T19:46:22-0500

First of all, let us start with the condition "p'(1)=3" and we will see why in a second :

Let's calculate "p'(x) = (x^2+ax+b)' = 2x+a" (because "(x^2)'=2x, (ax)'=a, b'=0" ). The condition "p'(1)=3" gives us "2\times 1 + a = 3" , equation where b vanishes and so we have a simple solution "a=1" .

Now we can proceed to the condition "p(1)=-6" , it gives us "1^2+a\times1+b=-6" , knowing that a=1, so we have "b= -8" .

So our final answer is: "p(x) = x^2+x-8" .